OpenStudy (ebonyserrano):

Three consecutive even integers have a sum of 48. Find the integers.

likeabossssssss (likeabossssssss):

what do u think it is?

OpenStudy (ebonyserrano):

I got 6,7,8

OpenStudy (ebonyserrano):

@likeabossssssss thanks for the help

OpenStudy (ebonyserrano):

it says it's incorrect

OpenStudy (ebonyserrano):

it said it would be 14,16, and 18

563blackghost (563blackghost):

We know three consecutive numbers equal 48, due to the fact that the number are even we would go up by multiples of 2. \(\huge\bf{(n)+(n+2)+(n+4)=48}\) Combine Like terms. \(\huge\bf{3n+6=48}\) Subtract 6 from both sides. \(\huge\bf{3n=42}\) Divide by 3. \(\huge\bf{n=14}\) Now input 14. \(\huge\bf{(14)=14}\) \(\huge\bf{14+2=16}\) \(\huge\bf{14+4=18}\) \(\huge\bf\color{red}{14,16,18}\) That should be your answer @Ebonyserrano

563blackghost (563blackghost):

At least I think thats how you do it...do you agree @KJSaif ?